OpenStudy (marigirl):

Solve the equation \frac{ (x^2+3x+2)(x^2-6x+8) }{ x^2-4 }=3x-12 When I go to solve it, I end up with (x+1)(x-4)=3x-12 (x-4)(x-2)=0 And x=4,2 BUT I know 2 is not a valid solution. How does this question work algebraically?

OpenStudy (marigirl):

\[\frac{ (x^2+3x+2)(x^2-6x+8) }{ x^2-4 }=3x-12\]

OpenStudy (solomonzelman):

\(\color{purple}{\displaystyle\frac{ (x^2+3x+2)(x^2-6x+8) }{ x^2-4 }=3x-12 }\)\(\tiny \\[0.6em]\) ------------------------------------------\(\tiny \\[0.6em]\) Factor: \(\color{purple}{\displaystyle\frac{ (x+1)(x+2)(x-2)(x-4) }{ (x-2)(x+2) }=3x-12 }\) your denominator can't be equal to 0, so assume \(\color{black}{\displaystyle x\ne\pm2}\). Then, since \(\color{black}{\displaystyle x\ne\pm2}\), you can divide top and bottom by \(\color{black}{\displaystyle (x-2)(x+2)}\), (because \(\color{black}{\displaystyle (x-2)(x+2)\ne0}\), so you wouldn't be dividing by 0). So you get: \(\color{purple}{\displaystyle (x+1)(x-4)=3x-12 }\) \(\color{purple}{\displaystyle (x+1)(x-4)=3(x-4) }\) Notice that if \(\color{black}{\displaystyle x-4=0 }\) your equation is satisfied, so \(\color{black}{\displaystyle x=4 }\) is therefore a solution (solution #1).

OpenStudy (solomonzelman):

So, having excluded the solution \(\color{black}{\displaystyle x=4 }\), we will assume \(\color{black}{\displaystyle x\ne4 }\), and then division by \(\color{black}{\displaystyle x-4 }\) on both sides is valid. This division gives you, \(\color{black}{\displaystyle x+1=3 }\) \(\color{black}{\displaystyle x=2 }\)

OpenStudy (solomonzelman):

But, since \(\color{black}{x=2}\) has been already excluded (that is your restriction for the denominator \(\color{black}{x^2-4}\) not to be \(\color{black}{0}\)), therefore the only working solution is \(\color{black}{x=4}\).